# Senior Member, IEEE

### v3

@@ 1 INTRODUCTION
Due to these mentioned reasons, recent advancements focus on reducing privacy concerns by using non-visual signal-based sensing methods [5, 6]. These signals contain no human facial information or visual photographs, thus protecting human privacy in IoT application scenarios [7, 8]. Millimeter-wave (mmWave) radar has emerged as a prominent technology due to its high resolution and cost-effectiveness. MmWave radar can accurately estimate human skeletal poses without capturing identifiable visual information, making it ideal for preserving privacy. This paper focuses on mmWave radar-based human skeletal pose estimation, a critical component in smart sensor systems for medical care and smart home IoT applications.
MmWave radar, a robust yet low-cost sensor, has been applied for environment perception in advanced driver-assistance systems (ADAS) [9, 10, 11, 12, 13, 14] and cooperative intelligent transportation systems (C-ITS) [15] in recent years. Considerable frontier research on mmWave radar has focused on achieving human activity sensing tasks. However, significant challenges remain. In the classical radar processing pipeline, various levels of data representation exist. For human skeletal pose estimation, point clouds and heatmaps are the most widely used forms of data representation. However, point cloud generation relies on hand-crafted parameters, and the process itself loses Doppler information and some positional data, leading to difficulties in capturing finer joint features. Heatmaps are often generated using the Fast Fourier Transform (FFT) method [16], which can extract range, Doppler, and angle of arrival (AoA) information from the raw radar signal to create different types of heatmaps. However, much of the information in the heatmap is considered redundant and not valued. Recently, many pose estimation works have been realized through a point cloud $+$ neural network pipeline[17] or a heatmap $+$ neural network pipeline. Some studies have focused on improving the deep learning capability of networks and algorithms to address the mmWave radar-based precision problem. However, the importance of capturing the information present in the radar signal itself is often neglected, leading to the loss of significant effective information.
@@ 2 RELATED WORK
2.1 Millimeter-wave Radar Indoor Sensing
The field of millimeter-wave (mmWave) radar for indoor sensing has gained substantial advancements in recent years, driven by the growing demand for non-invasive, privacy-preserving technologies for human presence and activity detection. MmWave radar offers unique advantages for indoor sensing applications. The mmWave frequency can penetrate non-metallic objects while providing high-resolution imaging capabilities [19]. Its utility is increasingly evident in through-wall sensing [20], human detection and recognition [21], and vital sign monitoring [22, 23, 24, 25].
Initially, researchers used radar signal processing to extract waveform features, combining them with target distance, orientation, altitude, and motion information to perform simple action classification[26]. Later, advanced technologies, e.g., deep learning, were introduced for more accurate indoor sensing projects based on basic mmWave radar data. For instance, Wang et al. proposed a sequence-to-sequence (seq2seq) 3D temporal convolutional network with a self-attention method to estimate human hand pose [27]. Yi et al. utilized a multi-person detection model based on long short-term memory (LSTM) to determine human presence and localize their positions using a single 60GHz mmWave radio [28]. Ding et al. developed a novel hybrid neural network model that explores multi-domain fusion of radar information by utilizing three kinds of 2D domain spectra. They combine 1D and 2D convolutional neural network and a recurrent neural network to capture rich features through multiple domains. [29] Wu et al. pioneered the work of segmenting human silhouettes from millimeter-wave RF signals by locating human reflections in radar frames and extracting features from surrounding signals with a human detection module and from aggregated frames with an attention-based mask generation module [30].
@@ 2.2 MmWave based Human Skeletal Pose Estimation
RF-Pose [31] is one of the pioneering works that first considered human skeleton reconstruction. Zhao et al. proposed a deep network from Frequency Modulated Continuous Wave(FMCW) radar signals to estimate coarse parts of the human skeleton under the supervision of visual information. They subsequently improved the model to successfully predict 14 human joints, including the head, neck, shoulders, elbows, wrists, hips, knees, and feet. RF-Pose and its follow-up work utilized expensive customized hardware systems. Nowadays, more studies are based on low-cost mass-produced industrial millimeter-wave radar and have demonstrated good performance. Ding et al. presented a kinematic-constrained learning architecture that incorporates kinematic constraints with neural network learning for skeleton estimation based on range-Doppler heatmaps [32]. Kong et al. performed a convolution operation on the Range-Doppler Profile to detect the corresponding ranges and designed a two-stream deep learning architecture to extract body shape and motion features to predict skeleton joint coordinates and reconstruct body posture [33]. Cao et al. developed a model that incorporates part-level range-Doppler maps for individual body parts with local kinematic constraints and global constraints for reconstructing the human skeleton [34]. Cao et al. proposed a two-stage framework for task-specific feature purification, which initially employs an adversarial auto-encoder to extract interpersonal-independent features, followed by a feature disentanglement module to eliminate components irrelevant to the pose [35].
@@ 3.1 FFT Branch: Radar Data Processing and Range-Doppler-Angle Map Generating
×$\times e^{-j2\pi\left(\frac{hn}{N}+\frac{im}{M}+\frac{jp}{P}+\frac{kq}{Q}\right)},$
where ${f(n,m,p,q)}$ represents the original input signal across ADC samples ${n}$, chirps ${m}$, horizontal antennas ${p}$, and vertical antennas ${q}$, and ${F(h,i,j,k)}$ denotes the results of the Fourier transform. ${N,M,P,Q}$ are the sizes of the dimensions, while ${h,i,j,k}$ are the indices in the frequency domain for the corresponding dimensions. This process yields detailed range-Doppler-azimuth-elevation maps. Additionally, a certain amount of chirps within a specific velocity range is uniformly sampled in the Doppler dimension to filter out irrelevant information in the radar signal. This strategy effectively eliminates extraneous signal components, thereby streamlining the dataset for enhanced processing efficiency.
@@ 3.2 ProbPE Branch: Probability Map Generating and Positional Encoding
where ${T}$ represents the threshold for the cell, ${X_{i}}$ represents the amplitude values of the reference cells surrounding the cell under test, ${N}$ represents the number of reference cells considered, and ${\alpha}$ is a scaling factor determined by the desired false alarm rate. If the intensity of the target surpasses that of interference noise, as ${I_{\mathrm{target}}>I_{\mathrm{noise}}\times T}$, it indicates the presence of the target.
@@
where ${P_{\mathrm{rae}}(r,\theta,\phi)}$ represents the probability map indicating the likelihood of detecting a target at range ${r}$, azimuth ${\theta}$, and elevation ${\phi}$. ${\hat{V}\mathrm{ra}(r,\theta)}$ represents the normalized range-azimuth vector, and ${\hat{V}\mathrm{re}(r,\phi)}$ denotes the normalized range-elevation vector.
@@ 3.3 Multi-format Feature Fusion and Estimation Head
where ${H}$ and ${\hat{H}}$ represent the initial keypoint heatmaps and refined keypoint heatmaps, respectively, and ${G}$ represents the generated ground truth keypoint heatmaps based on Gaussian distribution. To illustrate, the binary cross-entropy loss $L_{\mathrm{bce}}(H,G)$ is defined by
@@ 4 EXPERIMENTS AND RESULTS
This section introduces the implementation and detailed performance of our proposed method. Ablation experiments are also constructed to compare the prediction performance of incomplete models.
@@ 4.1 Dataset and Evaluation
The HuPR dataset was chosen for our experiments [18]. This dataset acquires data from two identical radars. One radar sensor is rotated $90^{\circ}$ in the antenna plane with respect to the other, where one radar focuses on the horizontal plane and the other on the vertical plane. Unlike some radar datasets that provide processed data, e.g., point clouds [41, 42, 43], heatmaps [44], etc., which lose the purity of radar data, HuPR provides raw radar ADC signals. Table 1 demonstrates a comparison of the data formats provided by mainstream accessible mmWave radar-based human posture datasets. HuPR dataset the only dataset that provides radar signals that have not been manually processed. This allows for more possibilities for our data processing and feature extraction methods.
@@
$Dataset$ $Radar$ $RawSignal$ $PointCloud$ $Heatmap$
@@ 4.2 Implementation Details
$Layers$ AP
$Head$ $Elbow$ $Wrist$ $Total$
@@ 5 CONCLUSION
In this paper, mmWave radar was applied to estimate human skeletal poses. This is an important component of smart sensor systems used in medical care IoT applications, particularly in addressing privacy concerns. ProbRadarM3F is designed to apply millimeter-wave radar for human skeletal pose estimation. This model significantly improves the extraction and use of hidden information from radar signals, by introducing the ProbPE branch, which generates probability maps based on estimating the likelihood of a specific position. These probability maps are then used to effectively extract position features using a positional encoding method. Additionally, we combined positional features with features obtained from the FFT branch to enhance the human body keypoint features produced by the model. Our experiments on the HuPR dataset demonstrated that ProbRadarM3F outperforms existing methods, indicating the effectiveness of our approach. Our results highlight that there is valuable information, such as positional data, present in the radar heatmap or even in the radar signal itself, which should not be considered redundant in human pose estimation tasks.

### v5

@@ 1 INTRODUCTION
Due to these mentioned reasons, recent advancements focus on reducing privacy concerns by using non-visual signal-based sensing methods [5, 6]. These signals contain no human facial information or visual photographs, thus protecting human privacy in IoT application scenarios [7, 8]. Millimeter wave (mmWave) radar has emerged as a prominent technology due to its high resolution and cost-effectiveness. MmWave radar can accurately estimate human skeletal poses without capturing identifiable visual information, making it ideal for preserving privacy. This paper focuses on mmWave radar-based human skeletal pose estimation, a critical component in smart sensor systems for medical care and smart home IoT applications.
MmWave radar, a robust yet low-cost sensor, has been applied for environment perception in advanced driver-assistance systems (ADAS) [9, 10, 11, 12, 13, 14] and cooperative intelligent transportation systems (C-ITS) [15] in recent years. Considerable frontier research on mmWave radar has focused on achieving human activity sensing tasks. However, significant challenges remain. In the classical radar processing pipeline, various levels of data representation exist. For human skeletal pose estimation, point clouds and heatmaps are the most widely used forms of data representation. However, point cloud generation relies on hand-crafted parameters, and the process itself loses Doppler information and some positional data, leading to difficulties in capturing finer joint features. Heatmaps are often generated using the Fast Fourier Transform (FFT) method [16], which can extract range, Doppler, and angle of arrival (AoA) information from the raw radar signal to create different types of heatmaps. However, much of the information in the heatmap is considered redundant and not valued. Recently, many pose estimation works have been realized through a point cloud ++ neural network pipeline[17] or a heatmap ++ neural network pipeline. Some studies have focused on improving the deep learning capability of networks and algorithms to address the mmWave radar-based precision problem. However, the importance of capturing the information present in the radar signal itself is often neglected, leading to the loss of significant effective information.
@@ 2 RELATED WORK
2.1 Millimeter wave Radar Indoor Sensing
The field of millimeter wave (mmWave) radar for indoor sensing has gained substantial advancements in recent years, driven by the growing demand for non-invasive, privacy-preserving technologies for human presence and activity detection. MmWave radar offers unique advantages for indoor sensing applications. The mmWave frequency can penetrate non-metallic objects while providing high-resolution imaging capabilities [19]. Its utility is increasingly evident in through-wall sensing [20], human detection and recognition [21], and vital sign monitoring [22, 23, 24, 25].
Initially, researchers used radar signal processing to extract waveform features, combining them with target distance, orientation, altitude, and motion information to perform simple action classification[26]. Later, advanced technologies, e.g., deep learning, were introduced for more accurate indoor sensing projects based on basic mmWave radar data. For instance, Wang et al. proposed a sequence-to-sequence (seq2seq) 3D temporal convolutional network with a self-attention method to estimate human hand pose [27]. Yi et al. utilized a multi-person detection model based on long short-term memory (LSTM) to determine human presence and localize their positions using a single 60GHz mmWave radio [28]. Ding et al. developed a novel hybrid neural network model that explores multi-domain fusion of radar information by utilizing three kinds of 2D domain spectra. They combine 1D and 2D convolutional neural network and a recurrent neural network to capture rich features through multiple domains [29]. Wu et al. pioneered the work of segmenting human silhouettes from millimeter wave RF signals by locating human reflections in radar frames and extracting features from surrounding signals with a human detection module and from aggregated frames with an attention-based mask generation module [30].
@@ 2.2 MmWave based Human Skeletal Pose Estimation
RF-Pose [31] is one of the pioneering works that first considered human skeleton reconstruction. Zhao et al. proposed a deep network from Frequency Modulated Continuous Wave(FMCW) radar signals to estimate coarse parts of the human skeleton under the supervision of visual information. They subsequently improved the model to successfully predict 14 human joints, including the head, neck, shoulders, elbows, wrists, hips, knees, and feet. RF-Pose and its follow-up work utilized expensive customized hardware systems. Nowadays, more studies are based on low-cost mass-produced industrial millimeter wave radar and have demonstrated good performance. Ding et al. presented a kinematic-constrained learning architecture that incorporates kinematic constraints with neural network learning for skeleton estimation based on range-Doppler heatmaps [32]. Kong et al. performed a convolution operation on the Range-Doppler Profile to detect the corresponding ranges and designed a two-stream deep learning architecture to extract body shape and motion features to predict skeleton joint coordinates and reconstruct body posture [33]. Cao et al. developed a model that incorporates part-level range-Doppler maps for individual body parts with local kinematic constraints and global constraints for reconstructing the human skeleton [34]. Cao et al. proposed a two-stage framework for task-specific feature purification, which initially employs an adversarial auto-encoder to extract interpersonal-independent features, followed by a feature disentanglement module to eliminate components irrelevant to the pose [35].
@@ 3.1 FFT Branch: Radar Data Processing and Range-Doppler-Angle Map Generating
$\times e^{-j2\pi\left(\frac{hn}{N}+\frac{im}{M}+\frac{jp}{P}+\frac{kq}{Q}\right)},$
where f(n,m,p,q){f(n,m,p,q)} represents the original input signal across ADC samples n{n}, chirps m{m}, horizontal antennas p{p}, and vertical antennas q{q}, and F(h,i,j,k){F(h,i,j,k)} denotes the results of the Fourier transform. N,M,P,Q{N,M,P,Q} are the sizes of the dimensions, while h,i,j,k{h,i,j,k} are the indices in the frequency domain for the corresponding dimensions. This process yields detailed range-Doppler-azimuth-elevation maps. Additionally, a certain amount of chirps within a specific velocity range is uniformly sampled in the Doppler dimension to filter out irrelevant information in the radar signal. This strategy effectively eliminates extraneous signal components, thereby streamlining the dataset for enhanced processing efficiency.
@@ 3.2 ProbPE Branch: Probability Map Generating and Positional Encoding
where T{T} represents the threshold for the cell, $X_{i}$} represents the amplitude values of the reference cells surrounding the cell under test, N{N} represents the number of reference cells considered, and $\alpha$} is a scaling factor determined by the desired false alarm rate. If the intensity of the target surpasses that of interference noise, as $I_{\mathrm{target}}>I_{\mathrm{noise}}\times T$}, it indicates the presence of the target.
@@
where $P_{\mathrm{rae}}(r,\theta,\phi)$} represents the probability map indicating the likelihood of detecting a target at range r{r}, azimuth $\theta$}, and elevation $\phi$}. $\hat{V}\mathrm{ra}(r,\theta)$} represents the normalized range-azimuth vector, and $\hat{V}\mathrm{re}(r,\phi)$} denotes the normalized range-elevation vector.
@@ 3.3 Multi-format Feature Fusion and Estimation Head
where H{H} and $^{\hat{H}}$ represent the initial keypoint heatmaps and refined keypoint heatmaps, respectively, and G{G} represents the generated ground truth keypoint heatmaps based on Gaussian distribution. To illustrate, the binary cross-entropy loss $L_{\mathrm{bce}}(H,G)$ is defined by
@@ 4 EXPERIMENTS AND RESULTS
This section introduces the implementation and detailed performance of our proposed method. Ablation experiments are also conducted to compare the prediction performance of incomplete models.
@@ 4.1 Dataset and Evaluation
The HuPR dataset was chosen for our experiments [18]. This dataset acquires data from two identical radars. One radar sensor is rotated $90^{\circ}$ in the antenna plane with respect to the other, with one radar focusing on the horizontal plane and the other on the vertical plane. Unlike some radar datasets that provide processed data, e.g., point clouds [41, 42, 43], heatmaps [44], etc., which lose the purity of radar data, HuPR provides raw radar ADC signals. Table 1 demonstrates a comparison of the data formats provided by mainstream accessible mmWave radar-based human posture datasets. The HuPR dataset is the only dataset that provides radar signals that have not been manually processed. This allows for more possibilities for our data processing and feature extraction methods.
@@
DatasetDataset RadarRadar RawSignalRawSignal PointCloudPointCloud HeatmapHeatmap
@@ 4.2 Implementation Details
LayersLayers AP
HeadHead ElbowElbow WristWrist TotalTotal
@@ 4.4 Ablation Study
  • $\bullet$
    To determine the impact of the ProbPE branch, the network was operated with only the FFT branch, excluding the integration of features from the positional encoding guided by probability maps. The experimental results are shown in Table 5.
@@ 5 CONCLUSION
In this paper, mmWave radar was applied to estimate human skeletal poses. This is an important component of smart sensor systems used in medical care IoT applications, particularly in addressing privacy concerns. ProbRadarM3F is designed to apply millimeter wave radar for human skeletal pose estimation. This model significantly improves the extraction and use of hidden information from radar signals, by introducing the ProbPE branch, which generates probability maps based on estimating the likelihood of a specific position. These probability maps are then used to effectively extract position features using a positional encoding method. Additionally, we combined positional features with features obtained from the FFT branch to enhance the human body keypoint features produced by the model. Our experiments on the HuPR dataset demonstrated that ProbRadarM3F outperforms existing methods, indicating the effectiveness of our approach. Our results highlight that there is valuable information, such as positional data, present in the radar heatmap or even in the radar signal itself, which should not be considered redundant in human pose estimation tasks.